CALCULATION OF VIBRATIONAL ENERGY LEVELS: VPT2 VS. A VARIATIONAL APPROACH \underline{M. E. HARDING}, J. VÁZQUEZ, J. F. STANTON, Institute for Theoretical Chemistry, Department of Chemistry and Biochemistry, University of Texas at Austin, Austin, TX 78712 , USA.

For several potentials a detailed comparison of vibrational energy levels obtained via variational calculations employing the complete vibrational Hamiltonian^{a} and second-order vibrational perturbation theory (VPT2) is carried out.

aJ. K. G. Watson, Mol. Phys., 15, 479 (1968):19, 465 (1970).